# AVA-Bench: Atomic Visual Ability Benchmark for Vision Foundation Models

### v1

@@ Abstract
The rise of vision foundation models (VFMs) calls for systematic evaluation. A common approach pairs VFMs with large language models (LLMs) as general-purpose heads, followed by evaluation on broad Visual Question Answering (VQA) benchmarks. However, this protocol has two key blind spots: (i) the instruction tuning data may not align with VQA test distributions, meaning a wrong prediction can stem from such data mismatch rather than a VFM’ visual shortcomings; (ii) VQA benchmarks often require multiple visual abilities, making it hard to tell whether errors stem from lacking all required abilities or just a single critical one. To address these gaps, we introduce AVA-Bench, the first benchmark that explicitly disentangles 14 Atomic Visual Abilities (AVAs)—foundational skills like localization, depth estimation, and spatial understanding that collectively support complex visual reasoning tasks. By decoupling AVAs and matching training and test distributions within each, AVA-Bench pinpoints exactly where a VFM excels or falters. Applying AVA-Bench to leading VFMs thus reveals distinctive “ability fingerprints,” turning VFM selection from educated guesswork into principled engineering. Notably, we find that a 0.5B LLM yields similar VFM rankings as a 7B LLM while cutting GPU hours by 8$\times$, enabling more efficient evaluation. By offering a comprehensive and transparent benchmark, we hope AVA-Bench lays the foundation for the next generation of VFMs.
@@ 1 Introduction
Vision Foundation Models (VFMs), pre-trained on large and diverse datasets, have become central to AI by providing transferable features for a wide range of downstream tasks [40, 2, 13, 5]. The variety of pre-training objectives and supervision signals has led to a proliferation of specialized VFMs—such as DINOv2 [65], CLIP [67], and SAM [43]—each excelling in distinct visual capabilities while often exhibiting interesting emergent properties [24, 6, 64]. Consequently, establishing a systematic and effective evaluation protocol for VFMs has become increasingly crucial.
Existing evaluation protocols can generally be categorized into two groups. The first focuses on task-specific capabilities, typically attaching tailored heads to VFMs, followed by fine-tuning and evaluation on dedicated datasets such as ImageNet for classification [30, 57] and COCO for detection or segmentation [79, 3]. To better capture the diverse and complex perception challenges of the real world, recent studies advocate a more generic approach that leverages large language models (LLMs) as general-purpose heads, evaluating VFMs on broad Visual Question Answering (VQA) benchmarks [54, 110, 12].
@@
We systematically benchmark leading VFMs trained under diverse objectives and data (see Figure 1), covering language-supervised (e.g., SigLIP-1/2 [82, 101], CLIP [67], InternVL-2.5 [9]), multimodal autoregressive (AIMv2 [20]), segmentation-supervised (SAM [43]), depth-supervised (MiDas [68]), contrastive self-supervised (DINOv2 [65]), and agglomerative models (RADIO [70]). We follow the standard protocol of adding an LLM on top, but fine-tune it separately for each AVA.
@@
Alongside our main study, we identify a more resource-efficient evaluation strategy. Existing LLM-based VFM evaluations typically rely on heavyweight models such as Vicuna-1.5(7B/13B) [80, 35], aiming for high absolute accuracy but incurring significant computational overhead. However, when the goal is to compare VFMs, we advocate prioritizing relative performance over absolute accuracy. We demonstrate that a lightweight 0.5B LLM preserves reliable VFM rankings while reducing evaluation costs by 8$\times$, making large-scale analysis substantially more practical. We also contextualize our findings within related work [24, 80] to offer a holistic understanding of existing VFMs and highlight directions for future research (see section 6).
@@ 2.1 Vision Foundation Models (VFMs)
A wide range of architectures and learning strategies have been explored to develop vision foundation models (VFMs). In general, they fall into two categories based on the nature of their training data. On one hand, VFMs are pre-trained purely on visual data. For instance, the Vision Transformer (ViT)  [16] is trained on labeled images using the supervised signal to capture visual representations. DINOv2  [65], in contrast, adopts a self-supervised learning approach, enabling the model to learn those features without labeling. Some VFMs are designed for specific vision tasks—for example, SAM [43] specializes in open-vocabulary segmentation, while MiDaS [68] focuses on monocular depth estimation. RADIO  [70] instead introduces a multi-teacher distillation framework that unifies the strengths of different VFMs (e.g., CLIP, DINOv2, and SAM) into a single efficient student model.
Another family of VFMs leverages image-text pairs. CLIP  [67] uses contrastive learning to align image and textual descriptions, while SigLIP  [82, 101] replaces the contrastive loss with a sigmoid one for more efficient training. Unlike the conventional language-guided VFMs that process images and text separately, AIMv2  [20] integrates visual and textual understanding into a single auto-regressive framework, which is simple yet effective. Due to their different architectures and learning objectives, VFMs may have varying strengths and limitations in visual understanding.
@@ 2.2 LLM-based Evaluation for VFMs
Unlike the traditional paradigm, where different perception tasks, such as classification, detection or segmentation, necessitate task-specific models [30, 2], the widespread adoption of Large Language Models (LLMs) as versatile interfaces has significantly shifted this paradigm [102, 91]. Reflecting this shift, recent studies advocate leveraging large language models (LLMs) as general-purpose heads and evaluating VFMs on broad Visual Question Answering (VQA) benchmarks [80]. Specifically, following the LLaVA approach, these studies utilize a two-stage framework: (i) pre-training a connector between frozen LLM and VFMs with image-text pairs for feature alignment (ii) fine-tuning both the connector and the LLM using instruction-tuning data, while keeping VFMs frozen during both stages (see Figure 5). This LLM-based evaluation protocol has rapidly gained popularity as it closely mirrors the contemporary multimodal LLM setting and effectively captures diverse real-world perception challenges.
@@ 3.1 Atomic Visual Abilities (AVAs)
The AVAs selected for AVA-Bench are grounded in a thorough literature analysis, including compositional text-to-images benchmarks [36, 41, 93] and VQA questions [25, 94, 1] (details in Appendix), focusing strictly on pure perceptual tasks and excluding non-perceptual reasoning skills (e.g., historical context, mathematical reasoning). Examples of each AVA can be found in Figure 3 and detailed definitions can be found in Appendix.
@@ 3.2 Dataset Curation
Constructing AVA-Bench required carefully isolating image–question pairs that specifically test individual AVAs. Existing MLLM and VQA datasets often blend multiple perceptual abilities, complicating direct assessments [81, 106, 100]. To achieve clear isolation, we assembled a comprehensive suite comprising image–question pairs from 26 diverse datasets, explicitly aligning each pair with a single targeted AVA. These datasets span a broad range of domains—including general scenes, wildlife (e.g., birds, fungi, plants), vehicles, indoor/outdoor settings, and remote-sensing imagery.
@@ 3.2.1 Localization
Localization is a crucial foundational AVA, essential for various VQA questions. We curate 34.8K localization-focused image–question pairs sourced from diverse datasets: Objects365 [73] and LVIS [29] (open-domain), iNaturalist-2021 [85] (birds and animals), and DIOR [49] (remote-sensing). To ensure clarity, we include images with a single instance of the target object and exclude objects with bounding box area smaller than 0.2% of the image.
@@ 3.2.2 Counting
We collect 11.3K image-question pairs from five datasets: VQAv2 [25] (open-domain), FSC [69] (open-domain counting), CARPK [33] (cars), Crowd Surveillance Dataset [48] (people), and LVIS [29] (open-domain). FSC, CARPK, and Crowd datasets explicitly cater to counting tasks, ensuring clear and distinct instance counts. LVIS instance masks are utilized to generate additional counting questions, complemented by counting-focused selections from VQAv2.
@@ 3.2.3 Spatial Reasoning
We curate 15.9K pairs from NYU-Depth V2 [74] for indoor scenes with object segmentation, and LVIS [29] and Objects365 [73] for diverse open-domain scenarios. For each image, we select two non-overlapping objects—one marked with a blue bounding box (the reference object) and another with a red bounding box (the target). The model is then asked to determine the relative spatial position of the target object with respect to the reference one, choosing from multiple-choice options: top-left, top-right, bottom-left, and bottom-right. We ensure that the two boxes do not overlap either horizontally or vertically to avoid ambiguity.
@@ 3.2.4 Visual Attribute
Orientation. We compile 10.1K pairs from two specialized datasets, CURE-OR [77] and EgoOrientBench [39], providing uncluttered images of objects from nine distinct orientations: front, back, left, right, top, front left, front right, back left, and back right.
Color. We curate 14K pair sourcing from from open-domain datasets, such as Objects365 [73] and LVIS [29]. To isolate the color recognition from object semantics or background clutter, we provide bounding boxes that tightly localize regions with minimal color variance, ensuring that the model focuses on RGB prediction.
Texture. We collect 13.2K pairs from four open-domain datasets: DTD [14], Kylberg [45], KTH-TIPS [78] and KTH-TIPS2 [60] with diverse texture types, such as striped, aluminum foil, and zigzagged.
Emotion. We gather 13.9K pairs from two publicly available datasets, RAF-DB [50] and ExpW [52], covering 7 annotated human emotional states: surprise, neutral, disgust, fear, happy, sad, and angry.
@@ 3.2.5 Depth Estimation
We curate depth-focused image-question pairs from NYU-Depth V2 [74] for indoor scenes and KITTI [23] for outdoor scenes.
@@ 3.2.6 Recognition
Action. We construct 15K pairs from the Moments in Time [63], a video dataset with diverse human actions. We extract the middle frame from each 3-second video for 302 distinct actions
Fine-grained. We curate 9K pairs from three diverse datasets: CUB-200-2011 [87] (birds), iNaturalist-2021 [85] (fungi, plants, animals), and FGVC-Aircraft [59] (objects). We randomly select 50 species across fungi, plants, and animals from iNaturalist [85] and 100 bird species from CUB, resulting in a total of 300 fine-grained classes.
Object. We curate 34.8K pairs from 4 datasets with diverse domains across 70 unique objects: Objects365 [73] and LVIS [29] (open-domain), iNaturalist-2021 [85] (birds and animals), and DIOR [49] (remote sensing). For each image, we provide a bounding box to eliminate the need for localization.
Scene. We curate 13.9K pairs from two datasets: Places434 [109] (open-domain) and AID [95] (remote sensing). The model is required to select the correct scene class from 30 randomly sampled options. The full set includes 464 classes spanning a wide range of scenes.
@@ 4 Evaluation Pipeline of AVA-Bench
To evaluate VFMs using AVA-Bench, we adopt the established LLM-based VFM evaluation protocol, employing the standard LLaVA-style two-stage training procedure to prepare the connector and LLM for VFM evaluation (details in subsection 2.2 and Figure 5). For each AVA in AVA-Bench, we fine-tune the connector and LLM while keeping the VFM frozen. Given the modest size of the training sets per AVA (typically around 6K–10K), we employ Parameter-Efficient Fine-Tuning (PEFT) [58, 32, 84], specifically Low-Rank Adaptation (LoRA) [34], to mitigate potential overfitting. Subsequently, the fine-tuned model is evaluated on the corresponding AVA-specific test sets.
@@ 4.1 Is a Heavyweight LLM Evaluator Necessary?
As discussed in subsection 2.2, traditional LLM-based VFM evaluations, following the LLaVA protocol, predominantly rely on heavyweight models such as Vicuna-1.5 (7B/13B)[54]., aiming for high absolute accuracy but incurring considerable computational costs. Nevertheless, a heavyweight LLM may not be mandatory for reliable comparative evaluations. When the goal is to compare VFMs, we advocate prioritizing relative performance over absolute metrics. As depicted in Figure 8, a significantly smaller 0.5B LLM (Qwen2) achieves similar relative VFM rankings comparable to a 7B Vicuna-1.5, while dramatically reducing evaluation costs by approximately 8$\times$ (additional details in the Appendix), making large-scale analysis substantially more practical. Thus, we utilize the lightweight 0.5B LLM for all subsequent experiments.
Figure 8: A much smaller 0.5B LLM achieves similar relative VFM rankings comparable to a 7B LLM, while dramatically reducing evaluation costs by approximately 8$\times$.
@@ 5.1 Setup
Metrics for AVAs. For absolute depth and counting AVAs, we utilized a normalized mean absolute error (MAE) relative to the ground-truth. This normalization ensures that errors involving greater distances or counts, which are inherently more challenging, are proportionally penalized less severely. Localization performance is evaluated using Generalized Intersection-over-Union (GIoU [71]), color via CIEDE2000 [56], and OCR through Average Normalized Levenshtein Similarity (ANLS) [4]. All other AVAs employ standard accuracy metrics. Please refer to Appendix for more metric details.
@@ 6 Discussion
Struggles of Non-Language-Aligned VFMs. Figure 7 and Figure 6 indicate that non-language-aligned VFMs, despite exhibiting strengths in certain vision-centric or low-to-mid-level AVAs (e.g., DINOv2 excelling in orientation and SAM leadning in color), typically underperform in most AVAs. To investigate the underlying cause of these shortcomings, we conducted a preliminary analysis focusing on the visual feature representations before and after the LLM connector. Using linear probing on max-pooled visual features from DINOv2, we observed a substantial accuracy degradation from 66.3% (pre-connector) to 25.67% (post-connector) for the fine-grained recognition AVA, aligning with findings reported in previous work [42]. This sharp performance drop suggests that critical visual information is often compromised during modality alignment processes. Recent studies have shown that partially unfreezing and fine-tuning the last few layers of VFMs can enhance performance [8]; however, this approach risks eroding the generalizability that initially makes these VFMs valuable. Alternatively, agglomerative models [70, 31] such as RADIO-2.1 demonstrate relatively robust performance on AVA-Bench, suggesting potential in combining specialized VFMs. Nevertheless, effectively aligning non-language-aligned VFMs to language modalities without sacrificing their inherent visual strengths remains a challenge, highlighting an important avenue for future research.
@@
Platonic Representation Hypothesis Holds? Recent research suggests that large-scale training may lead VFMs toward converging onto similar representational spaces–Platonic Representation Hypothesis [38]. Our findings partially support this hypothesis: for certain AVAs such as object recognition, texture, and relative depth estimation, VFMs demonstrate universally strong performance, indicative of similar underlying visual representations irrespective of training objectives. Conversely, significant performance disparities among VFMs in most other AVAs, as revealed by our evaluation, indicate limitations to the generality of this hypothesis. Thus, our study suggests that the Platonic Representation Hypothesis might only hold in certain cases and warrants further empirical scrutiny in more diverse and challenging contexts.
@@
What Are Things Going From Here? While MLLM have demonstrated remarkable versatility, they are not universally effective in all scenarios, especially in specialized domains [10, 53]. Thus, there is a growing necessity for developing specialized MLLMs [44]. Currently, selecting appropriate VFMs for such customized MLLMs remains largely heuristic [80, 75, 51]. Our work provides actionable insights that transform this selection process from heuristic guesswork into principled engineering. By clearly identifying AVA-specific strengths and weaknesses, practitioners can now systematically choose VFMs to precisely address the particular visual demands of targeted downstream tasks. Moreover, AVA-Bench represents a critical step towards developing next-generation VFMs by providing a systematic, diagnostic, and comprehensive evaluation framework. This benchmark enables VFM developers to accurately pinpoint specific deficiencies and implement targeted improvements, fostering the creation of more robust, versatile, and well-rounded VFMs in the future.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Vision Foundation Models (VFMs), pre-trained on large and diverse datasets, have become central to AI by providing transferable features for a wide range of downstream tasks khan2022transformers ; awais2025foundation ; chowdhury2025prompt ; bommasani2021opportunities . The variety of pre-training objectives and supervision signals has led to a proliferation of specialized VFMs—such as DINOv2 oquab2023dinov2 , CLIP radford2021learning , and SAM kirillov2023segment —each excelling in distinct visual capabilities while often exhibiting interesting emergent properties goldblum2023battle ; caron2021emerging ; naseer2021intriguing . Consequently, establishing a systematic and effective evaluation protocol for VFMs has become increasingly crucial.
Existing evaluation protocols can generally be categorized into two groups. The first focuses on task-specific capabilities, typically attaching tailored heads to VFMs, followed by fine-tuning and evaluation on dedicated datasets such as ImageNet for classification han2022survey ; mai2024fine and COCO for detection or segmentation thisanke2023semantic ; balachandran2024eureka . To better capture the diverse and complex perception challenges of the real world, recent studies advocate a more generic approach that leverages large language models (LLMs) as general-purpose heads, evaluating VFMs on broad Visual Question Answering (VQA) benchmarks liu2023visual ; zhu2023minigpt ; chowdhery2023palm .
@@
We systematically benchmark leading VFMs trained under diverse objectives and data (see Figure 1), covering language-supervised (e.g., SigLIP-1/2 tschannen2025siglip ; zhai2023sigmoid , CLIP radford2021learning , InternVL-2.5 chen2024expanding ), multimodal autoregressive (AIMv2 fini2024multimodal ), segmentation-supervised (SAM kirillov2023segment ), depth-supervised (MiDas ranftl2020towards ), contrastive self-supervised (DINOv2 oquab2023dinov2 ), and agglomerative models (RADIO ranzinger2024radio ). We follow the standard protocol of adding an LLM on top, but fine-tune it separately for each AVA.
@@
[…]
@@ 2.1 Vision Foundation Models (VFMs)
A wide range of architectures and learning strategies have been explored to develop vision foundation models (VFMs). In general, they fall into two categories based on the nature of their training data. On one hand, VFMs are pre-trained purely on visual data. For instance, the Vision Transformer (ViT)  dosovitskiy2020image is trained on labeled images using the supervised signal to capture visual representations. DINOv2  oquab2023dinov2 , in contrast, adopts a self-supervised learning approach, enabling the model to learn those features without labeling. Some VFMs are designed for specific vision tasks—for example, SAM kirillov2023segment specializes in open-vocabulary segmentation, while MiDaS ranftl2020towards focuses on monocular depth estimation. RADIO  ranzinger2024radio instead introduces a multi-teacher distillation framework that unifies the strengths of different VFMs (e.g., CLIP, DINOv2, and SAM) into a single efficient student model.
Another family of VFMs leverages image-text pairs. CLIP  radford2021learning uses contrastive learning to align image and textual descriptions, while SigLIP  tschannen2025siglip ; zhai2023sigmoid replaces the contrastive loss with a sigmoid one for more efficient training. Unlike the conventional language-guided VFMs that process images and text separately, AIMv2  fini2024multimodal integrates visual and textual understanding into a single auto-regressive framework, which is simple yet effective. Due to their different architectures and learning objectives, VFMs may have varying strengths and limitations in visual understanding.
@@ 2.2 LLM-based Evaluation for VFMs
Unlike the traditional paradigm, where different perception tasks, such as classification, detection or segmentation, necessitate task-specific models han2022survey ; awais2025foundation , the widespread adoption of Large Language Models (LLMs) as versatile interfaces has significantly shifted this paradigm zhang2024mm ; wu2023multimodal . Reflecting this shift, recent studies advocate leveraging large language models (LLMs) as general-purpose heads and evaluating VFMs on broad Visual Question Answering (VQA) benchmarks tong2024cambrian . Specifically, following the LLaVA approach, these studies utilize a two-stage framework: (i) pre-training a connector between frozen LLM and VFMs with image-text pairs for feature alignment (ii) fine-tuning both the connector and the LLM using instruction-tuning data, while keeping VFMs frozen during both stages (see Figure 5). This LLM-based evaluation protocol has rapidly gained popularity as it closely mirrors the contemporary multimodal LLM setting and effectively captures diverse real-world perception challenges.
@@ 3.1 Atomic Visual Abilities (AVAs)
The AVAs selected for AVA-Bench are grounded in a thorough literature analysis, including compositional text-to-images benchmarks huang2023t2i ; kil2024mllm ; wu2024conceptmix and VQA questions goyal2017making ; RealWorldQA ; ainslie2023gqa (details in Appendix), focusing strictly on pure perceptual tasks and excluding non-perceptual reasoning skills (e.g., historical context, mathematical reasoning). Examples of each AVA can be found in Figure 3 and detailed definitions can be found in Appendix.
@@ 3.2 Dataset Curation
Constructing AVA-Bench required carefully isolating image–question pairs that specifically test individual AVAs. Existing MLLM and VQA datasets often blend multiple perceptual abilities, complicating direct assessments tong2024eyes ; zhang2024mme ; yu2023mm . To achieve clear isolation, we assembled a comprehensive suite comprising image–question pairs from 26 diverse datasets, explicitly aligning each pair with a single targeted AVA. These datasets span a broad range of domains—including general scenes, wildlife (e.g., birds, fungi, plants), vehicles, indoor/outdoor settings, and remote-sensing imagery.
@@ 3.2.1 Localization
Localization is a crucial foundational AVA, essential for various VQA questions. We curate 34.8K localization-focused image–question pairs sourced from diverse datasets: Objects365 shao2019objects365 and LVIS gupta2019lvis (open-domain), iNaturalist-2021 van2021benchmarking (birds and animals), and DIOR li2020object (remote-sensing). To ensure clarity, we include images with a single instance of the target object and exclude objects with bounding box area smaller than 0.2% of the image.
@@ 3.2.2 Counting
We collect 11.3K image-question pairs from five datasets: VQAv2 goyal2017making (open-domain), FSC ranjan2021learning (open-domain counting), CARPK hsieh2017drone (cars), Crowd Surveillance Dataset li2022video (people), and LVIS gupta2019lvis (open-domain). FSC, CARPK, and Crowd datasets explicitly cater to counting tasks, ensuring clear and distinct instance counts. LVIS instance masks are utilized to generate additional counting questions, complemented by counting-focused selections from VQAv2.
@@ 3.2.3 Spatial Reasoning
We curate 15.9K pairs from NYU-Depth V2 silberman2012indoor for indoor scenes with object segmentation, and LVIS gupta2019lvis and Objects365 shao2019objects365 for diverse open-domain scenarios. For each image, we select two non-overlapping objects—one marked with a blue bounding box (the reference object) and another with a red bounding box (the target). The model is then asked to determine the relative spatial position of the target object with respect to the reference one, choosing from multiple-choice options: top-left, top-right, bottom-left, and bottom-right. We ensure that the two boxes do not overlap either horizontally or vertically to avoid ambiguity.
@@ 3.2.4 Visual Attribute
Orientation. We compile 10.1K pairs from two specialized datasets, CURE-OR Temel2018_ICMLA and EgoOrientBench jung2024right , providing uncluttered images of objects from nine distinct orientations: front, back, left, right, top, front left, front right, back left, and back right.
Color. We curate 14K pair sourcing from from open-domain datasets, such as Objects365 shao2019objects365 and LVIS gupta2019lvis . To isolate the color recognition from object semantics or background clutter, we provide bounding boxes that tightly localize regions with minimal color variance, ensuring that the model focuses on RGB prediction.
Texture. We collect 13.2K pairs from four open-domain datasets: DTD cimpoi14describing , Kylberg kylberg2011kylberg , KTH-TIPS kth-tips_dataset and KTH-TIPS2 article with diverse texture types, such as striped, aluminum foil, and zigzagged.
Emotion. We gather 13.9K pairs from two publicly available datasets, RAF-DB li2019reliable and ExpW lian2020expression , covering 7 annotated human emotional states: surprise, neutral, disgust, fear, happy, sad, and angry.
@@ 3.2.5 Depth Estimation
We curate depth-focused image-question pairs from NYU-Depth V2 silberman2012indoor for indoor scenes and KITTI geiger2013vision for outdoor scenes.
@@ 3.2.6 Recognition
Action. We construct 15K pairs from the Moments in Time monfort2019moments , a video dataset with diverse human actions. We extract the middle frame from each 3-second video for 302 distinct actions
Fine-grained. We curate 9K pairs from three diverse datasets: CUB-200-2011 wah2011caltech (birds), iNaturalist-2021 van2021benchmarking (fungi, plants, animals), and FGVC-Aircraft maji2013fine (objects). We randomly select 50 species across fungi, plants, and animals from iNaturalist van2021benchmarking and 100 bird species from CUB, resulting in a total of 300 fine-grained classes.
Object. We curate 34.8K pairs from 4 datasets with diverse domains across 70 unique objects: Objects365 shao2019objects365 and LVIS gupta2019lvis (open-domain), iNaturalist-2021 van2021benchmarking (birds and animals), and DIOR li2020object (remote sensing). For each image, we provide a bounding box to eliminate the need for localization.
Scene. We curate 13.9K pairs from two datasets: Places434 zhou2017places (open-domain) and AID xia2017aid (remote sensing). The model is required to select the correct scene class from 30 randomly sampled options. The full set includes 464 classes spanning a wide range of scenes.
@@ 4 Evaluation Pipeline of AVA-Bench
To evaluate VFMs using AVA-Bench, we adopt the established LLM-based VFM evaluation protocol, employing the standard LLaVA-style two-stage training procedure to prepare the connector and LLM for VFM evaluation (details in subsection 2.2 and Figure 5). For each AVA in AVA-Bench, we fine-tune the connector and LLM while keeping the VFM frozen. Given the modest size of the training sets per AVA (typically around 6K–10K), we employ Parameter-Efficient Fine-Tuning (PEFT) mai2025lessons ; houlsby2019parameter ; tu2023visual , specifically Low-Rank Adaptation (LoRA) hu2022lora , to mitigate potential overfitting. Subsequently, the fine-tuned model is evaluated on the corresponding AVA-specific test sets.
@@ 4.1 Is a Heavyweight LLM Evaluator Necessary?
[…]
@@ 5.1 Setup
Metrics for AVAs. For absolute depth and counting AVAs, we utilized a normalized mean absolute error (MAE) relative to the ground-truth. This normalization ensures that errors involving greater distances or counts, which are inherently more challenging, are proportionally penalized less severely. Localization performance is evaluated using Generalized Intersection-over-Union (GIoU rezatofighi2019generalized ), color via CIEDE2000 luo2001development , and OCR through Average Normalized Levenshtein Similarity (ANLS) biten2019scene . All other AVAs employ standard accuracy metrics. Please refer to Appendix for more metric details.
@@ 6 Discussion
Struggles of Non-Language-Aligned VFMs. Figure 7 and Figure 6 indicate that non-language-aligned VFMs, despite exhibiting strengths in certain vision-centric or low-to-mid-level AVAs (e.g., DINOv2 excelling in orientation and SAM leadning in color), typically underperform in most AVAs. To investigate the underlying cause of these shortcomings, we conducted a preliminary analysis focusing on the visual feature representations before and after the LLM connector. Using linear probing on max-pooled visual features from DINOv2, we observed a substantial accuracy degradation from 66.3% (pre-connector) to 25.67% (post-connector) for the fine-grained recognition AVA, aligning with findings reported in previous work finer . This sharp performance drop suggests that critical visual information is often compromised during modality alignment processes. Recent studies have shown that partially unfreezing and fine-tuning the last few layers of VFMs can enhance performance chen2024sharegpt4v ; however, this approach risks eroding the generalizability that initially makes these VFMs valuable. Alternatively, agglomerative models ranzinger2024radio ; heinrich2024radio such as RADIO-2.1 demonstrate relatively robust performance on AVA-Bench, suggesting potential in combining specialized VFMs. Nevertheless, effectively aligning non-language-aligned VFMs to language modalities without sacrificing their inherent visual strengths remains a challenge, highlighting an important avenue for future research.
@@
Platonic Representation Hypothesis Holds? Recent research suggests that large-scale training may lead VFMs toward converging onto similar representational spaces–Platonic Representation Hypothesis platonic . Our findings partially support this hypothesis: for certain AVAs such as object recognition, texture, and relative depth estimation, VFMs demonstrate universally strong performance, indicative of similar underlying visual representations irrespective of training objectives. Conversely, significant performance disparities among VFMs in most other AVAs, as revealed by our evaluation, indicate limitations to the generality of this hypothesis. Thus, our study suggests that the Platonic Representation Hypothesis might only hold in certain cases and warrants further empirical scrutiny in more diverse and challenging contexts.
@@
What Are Things Going From Here? While MLLM have demonstrated remarkable versatility, they are not universally effective in all scenarios, especially in specialized domains cheng2024domain ; liang2024comprehensive . Thus, there is a growing necessity for developing specialized MLLMs kumar2024diagnostics ; li2025visual . Currently, selecting appropriate VFMs for such customized MLLMs remains largely heuristic tong2024cambrian ; sun2025multi ; li2025openvision . Our work provides actionable insights that transform this selection process from heuristic guesswork into principled engineering. By clearly identifying AVA-specific strengths and weaknesses, practitioners can now systematically choose VFMs to precisely address the particular visual demands of targeted downstream tasks. Moreover, AVA-Bench represents a critical step towards developing next-generation VFMs by providing a systematic, diagnostic, and comprehensive evaluation framework. This benchmark enables VFM developers to accurately pinpoint specific deficiencies and implement targeted improvements, fostering the creation of more robust, versatile, and well-rounded VFMs in the future.